OpenStudy (coconut):

The basketball team scored 12 points the first game and 40 points the eighth game of the season. How many points will they earn with this pattern on the 14th game of the season?

OpenStudy (anonymous):

I'm not sure that defines a pattern?

OpenStudy (anonymous):

game 1 = 12 game 8 = a1 + d7 40 = 12 + d7 28 = d7 d = 4 game 14 = a1 + d*13 game 14 = 12 + 4*13 game 14 = 12 + 52 game 14 = 64

OpenStudy (anonymous):

between game 1 and 8, they increased 28 points. assuming an equal increase for each of the 7 games thru game 8, then they increased 28/7=4 pts per game. So we start with 12 and and for each numbered game (2nd, 3rd, etc) we add 4*(n-1). So the formula is 12+4(n-1). So the answer is 64

OpenStudy (anonymous):

I think that the quesiton given is incomplete. Two numbers don't define a pattern unless given some other information. You could come up with any crazy number pattern that includes those two in the right place...

OpenStudy (anonymous):

Scary, you are right. That is why I made the assumption that I did.

OpenStudy (anonymous):

I think the pattern is this: in odd games they score 12 points and in even games they score 40 points. So in the 14th game they'll score 40 points since it's even.

OpenStudy (coconut):

my options are 60 62 64 66

OpenStudy (anonymous):

Go with 64. The assumption of equal increase seems to have been a good one.

OpenStudy (anonymous):

64, then